OpenStudy (anonymous):

It is known Neptune is 30 AU from Sun, it takes 164 years to complete one orbit around the Sun. Use this info. alone to estimate the mass of the Sun. Should I use Newton's version of Kepler's 3rd law?

OpenStudy (vincent-lyon.fr):

Yes.

OpenStudy (anonymous):

But it says "using the info alone" so thats just orbital period, semi-major axis and unknown mass

OpenStudy (anonymous):

if i use newton's version then the G is an extra info

OpenStudy (vincent-lyon.fr):

You can't help it. G is a universal constant, so you are probably allowed to use it.

OpenStudy (anonymous):

Ok so can u tell me whats the problem of this equation? \[(164*365*24*3600)^{2}=\frac{(4\pi ^{2})[(30)(1.49*10^{8}])}{(6.67*10^{-11})(M)}\]

OpenStudy (vincent-lyon.fr):

I think you have forgotten to cube the radius of the orbit.

OpenStudy (anonymous):

oh right, i forgot to type in that but other than that?

OpenStudy (anonymous):

cus i can't get 1.99*10^30

OpenStudy (vincent-lyon.fr):

1au = 150 million km = 1.5 x 10¹¹ m

OpenStudy (anonymous):

dont i need to convert m to km? and even tho i used 10^11, i still can't get the ans.......